OpenStudy (anonymous):

cos^2(x)+cos(x)sin(x)=0 solve for x

myininaya (myininaya):

Try factoring

OpenStudy (anonymous):

I factored a cos(x) out but then im stuck

myininaya (myininaya):

cos(x)( ? + ? ) =0 So what should be in the places of those ?'s once you factor out cos(x)?

OpenStudy (anonymous):

cos(x)(cosx+sinx)

myininaya (myininaya):

Yes. Set both factors equal to 0. cos(x)=0 ; cos(x)+sin(x)=0

myininaya (myininaya):

Solve cos(x)=0 for x And find the value x so that cos(x) and sin(x) are opposites (i.e cos(x)=-sin(x))

OpenStudy (anonymous):

cosx=pie/2 ?

myininaya (myininaya):

cos(x)=0 for x=pi/2 cos(x) does not equal pi/2 But one value that satisfies cos(x)=0 is x=pi/2 Are there any other values?

OpenStudy (anonymous):

x=3pi/2

myininaya (myininaya):

Ok and you can also say we have x=pi/2+2npi where n is an integer or x=3pi/2+2n pi where n is an integer

myininaya (myininaya):

We also need to find the values of x where cos(x) and sin(x) have opposite values.

myininaya (myininaya):

 As you can see they will be opposite in quadrant 2 and 4.

myininaya (myininaya):

 These points is where cos(x) and sin(x) are opposite. Now what angle are these points at from the initial ray?